Hello...um i need to find the Value Of X and Y
4x=-5y+40
x=-y+9
i know itz pretty easy but i kinda forget everything during the summer...can u plz help me..it will b greatly appreciated =]

Answer by checkley77(12844)

4x=-5y+40
x=-y+9 MULTIPLY THIS EQUATION BY -4 7 ADD.
-4X=4Y-36
4X=-5Y+40
--------------
0=-Y+4
Y=4 ANSWER.
X=-4+9
X=5 ANSWER.
PROOF:
4*5=-5*4+40
20=-20+40
20=20